# Hallucination Detox: Sensitive Neuron Dropout (SeND) for Large Language Model Training

### v1

@@ Abstract
As large language models (LLMs) become increasingly deployed across various industries, concerns regarding their reliability, particularly due to hallucinations—outputs that are factually inaccurate or irrelevant to user input—have grown. Our research investigates the relationship between the training process and the emergence of hallucinations to address a key gap in existing research that focuses primarily on post hoc detection and mitigation strategies. Using models from the Pythia suite (70M–12B parameters) and several hallucination detection metrics, we analyze hallucination trends throughout training and explore LLM internal dynamics. We introduce SEnsitive Neuron Dropout (SeND), a novel training protocol designed to mitigate hallucinations by reducing variance during training. SeND achieves this by deterministically dropping neurons with significant variability on a dataset, referred to as Sensitive Neurons. In addition, we develop an unsupervised hallucination detection metric, Efficient EigenScore (EES), which approximates the traditional EigenScore in 2x speed. This efficient metric is integrated into our protocol, allowing SeND to be both computationally scalable and effective at reducing hallucinations. Our empirical evaluation demonstrates that our approach improves LLM reliability at test time by up to 40% compared to normal training while also providing an efficient method to improve factual accuracy when adapting LLMs to domains such as Wikipedia and Medical datasets.
@@ 1.1 Motivation
In the age of increasingly sophisticated and ever growing Large Language Models (LLMs), we see these models being utilized in every industry imaginable. However, concerns surrounding their reliability and safety have escalated due to their exploitation and the errors that arise when they are used by the general public. One of these concerning areas discovered by the scientific community is the phenomenon of hallucinations - LLMs producing content that may not align with real-world facts, the user’s input, or training data it has seen in the past (Huang et al., 2023a). In our research, we target a specific field of hallucinations called confabulations. Confabulations are a type of hallucination where the LLM generates different responses given the same or similar inputs. This can be harmful when the generations alter between correct and factually incorrect responses.
@@
To explore these hallucination trends, we analyze models ranging from 70 million to 12 billion parameters within EleutherAI’s Pythia suite (Biderman et al., 2023), assessing them across various training checkpoints and tasks. Our goal is to validate the oscillatory behavior observed in prior studies (Li et al., 2024) through evaluation metrics from HuggingFace and EleutherAI (Hong et al., 2024; Gao et al., 2024a), and to explore the correlation between model size, training progression, and hallucination patterns.
In response to the identified variance, we introduce a novel training protocol called Sensitive Neuron Dropout (SeND). SeND is designed to emphasize confident learning of facts, and in turn reduce the likelihood of confabulations, rather than solely minimizing the stochastic gradient descent (SGD) loss (e.g., cross-entropy). By selectively dropping Sensitive Neurons—those that exhibit significant fluctuations in contextual embeddings throughout training—SeND acts as a regularization technique that reduces hallucination variance and enhances the model’s factual certainty. This provides a more reliable criterion for determining training termination, ensuring models not only achieve loss convergence but also display stable factual confidence. To maintain efficiency as model size and inference count increase, we propose the Efficient EigenScore (EES), an approximation metric for hallucination detection. EES replaces EigenScore (Chen et al., 2024), the primary metric used in our experiments, offering a scalable solution with high correlation to the original EigenScore.
@@
Sensitive Neuron Dropout (SeND), a training-time method designed to reduce hallucination variance and increase model factual confidence during training.
@@ 1.2 Related Work
The majority of research on hallucinations in language models has focused on detecting and mitigating this phenomenon rather than explaining its underlying causes. Recent techniques can be categorized into two main approaches: those that rely on output text or model probabilities at inference time (Manakul et al., 2023; Joshi et al., 2017; Li et al., 2023) and those that utilize internal representations or hidden layers of the model (Su et al., 2024; Chen et al., 2024; Kossen et al., 2024). While the former has demonstrated effectiveness, the latter offers deeper insights but often comes with computational trade-offs. Additionally, methods like Reinforcement Learning with Human Feedback (RLHF) have gained traction for enhancing model reliability (Yu et al., 2024). However, many of these post-hoc solutions enhance factual accuracy by layering algorithms atop pre-trained models, which can be inefficient. Our work addresses this gap by focusing on the internal dynamics of the model that contribute to hallucinations.
We use several metrics for evaluation, including Halueval (Li et al., 2023), FactScore (Min et al., 2023), SelfCheckGPT (Manakul et al., 2023), and XSum (Narayan et al., 2018), to validate our findings across different tasks. Given that the internal dynamics of the model have proven to be reliable candidates for assessing certainty and hallucination likelihood, we leverage methodologies such as EigenScore (Chen et al., 2024) and Semantic Entropy (Kossen et al., 2024), which detect hallucination risk by analyzing the variability in multiple high-temperature outputs. In our experiments, we use the EigenScore metric alongside the HELM dataset created by Su et al. (2024) to detect hallucinations during training and in the development of SeND.
Regularization techniques have been introduced to fix the issue of variability, notably random neuron dropout, used to reduce the variance and ensure that no neuron is overpowering others (Srivastava et al., 2014; Baldi & Sadowski, 2013). Work such as that done by Santra et al. (2020); Ba & Frey (2013) aims to modify random neuron dropout to change the way neurons are dropped to a more deterministic, precise manner. This has allowed the authors to drop unimportant connections in a deep neural network to ensure that class discriminative information is propagated through the model correctly (Santra et al., 2020). Inspired by this, our aim is to target hallucinatory neurons in our models to ensure that factual information is propagated through.
A significant drawback of state-of-the-art detection techniques, particularly those relying on internal model dynamics, is their efficiency. Existing methods often necessitate multiple inferences and embedding generations, making the spectral analysis of embedding matrices computationally intensive and increasingly impractical as models and datasets grow (Chen et al., 2024; Su et al., 2024). To address these challenges, we propose the use of spectral theory for efficient approximation. This approach enables scalable hallucination detection while maintaining performance. By utilizing tools such as the Density of States (DOS) and the kernel polynomial method (KPM) for approximating EigenScore (Huang et al., 2023b; Lin et al., 2014), we aim to enhance the efficiency of our analysis in the context of confabulations, which we will demonstrate empirically with EES and SeND.
@@ 2 Oscillatory Behaviour Validation
The training epochs of a transformer model can be vital in understanding the dynamics of how the model learns, particularly when trained on an unsupervised loss with stochastic gradient. Beyond the model’s architecture and the data itself, numerous factors influence the learning process such as:
Whether the loss function penalizes the learner for factual mistakes it makes or if it primarily tries to force the model to memorize the data.
While our paper does not aim to address the broader debate on whether language models truly understand language or rely heavily on memorization, our analysis of training dynamics through multiple epochs shows that reducing stochastic gradient loss does not necessarily correspond to reducing hallucinations, verifying the results Li et al. (2024) showed for the oscillatory behaviour of LLMs in hallucination during training. This highlights the need for further investigation into the relationship between optimization and factual accuracy in LLMs.
2.1 Methods
In our study, we utilize Eleuther AI’s Pythia and LMEval tools (Biderman et al., 2023; Gao et al., 2024a), to examine the development and evolution of LLMs throughout the training process. Pythia comprises a suite of 16 LLMs, all trained on public data in the same sequential order, with sizes ranging from 70 million to 12 billion parameters, 8 of which we use for our experiments. We used 20 equally spaced training checkpoints from the start to the finish for our analysis. We chose Pythia as it is based on GPT-Neo X (Black et al., 2022), which shares a similar foundational architecture to other state-of-the-art language models. Pythia’s comprehensive package of models makes it particularly suitable for our analysis, allowing us to conduct a thorough examination of the development and evolution of LLMs throughout the training process.
These models are evaluated at each checkpoint on a variety of hallucination/fact-checking metrics. To do this, we leverage the HuggingFace Hallucination Leaderboard (Hong et al., 2024), which offers comprehensive benchmarks for our experiments. There are two main components to our evaluations: Summarization and Self-consistency. For summarization, models are evaluated under the XSum dataset Narayan et al. (2018) where the model is given a dataset of BBC news articles and must give summaries of each article. A higher Rouge1 score on XSum means the data is aligning better with the provided reference summary. SelfCheckGPT (Manakul et al., 2023) is used to see if the model is uncertain with respect to a dataset of prompts for self-consistency. When the SelfCheckGPT score is high, it means the model is more likely to contradict itself on the given input and therefore, more likely to hallucinate.
@@ 2.2 How do the established iterative training processes influence LLM hallucinations?
The analysis of hallucination oscillations, as shown in Figure 1, indicates a consistent pattern across different models: oscillations persist throughout training from the initial to the final checkpoint. This finding highlights the uncertainty of halting training solely based on the convergence of training loss to its minimum to minimize hallucination. For example, in the evaluation plot of XSUM with the 12B model (1(b)), the optimal value for the hallucination metric occurs at the 60000th checkpoint, despite the model’s unsupervised loss converging by step 140000, leading to the termination of training. This evidence challenges the notion that optimizing solely for unsupervised loss in SGD guarantees learning the most accurate representation of the data. This observation is seen more drastically in 1(a), where model size has no effect on the performance of SelfCheckGPT. Instead, we observe exaggerated oscillatory behaviour within self-consistency, meaning that model size is even less effective at tackling the issue of confabulations. One potential solution to mitigate this issue could involve incorporating a regularization term into the unsupervised loss based on a hallucination detection metric discussed in Section 4.
@@ 2.3 How does model complexity affect the emergence of hallucinations throughout training?
An analysis of hallucination detection metrics reveals a diminishing rate of improvement with increased model scaling, particularly up to the 12B parameter size (Figure 1(b)). This suggests that beyond a certain point, even though there is improvement in the hallucinations, larger models do not significantly reduce hallucinations, indicating that scaling alone is not sufficient for building robust models. Instead, more refined approaches are needed to address the underlying variability in model behavior. For the following experiments, we focus on the Pythia 1B model.
@@ 3.1 Sensitive Neurons
Where is the activation of one input , is the number of tokens in the sequence, and is the subtraction to get the penultimate layer index and the formula is adapted from Su et al. (2024). The penultimate layer of the LLM, being the layer closest to the output probabilities, is our primary focus for hallucination analysis due to its rich information about output certainty.
@@
Let denote the embedding of data point at neuron of the contextual embedding after checkpoint/epoch . Then we define the net change formula as
With these definitions, we can now describe the crux of our investigation: Sensitive Neurons. These Sensitive Neurons give us key parts of the model that we will prove contribute to the hallucination of LLM models. They can be used to adapt training procedures for lowering hallucination variation during training and better overall confidence at inference time. In essence, Sensitive Neurons are embedding indices in the sentence embedding from definition 3.1 that experience drastic changes between checkpoints/epochs of the training, something we believe is related to the oscillatory behaviour in hallucination performance. When finding the most Sensitive Neurons, we typically want to select the top neurons for a specific data point’s representation. In our investigation we set .
Definition 3.3 (Sensitive Neurons).
@@
where is the total variability during the last checkpoints and the most Sensitive Neurons are
@@
The above definition of Sensitive Neurons is then applied to LLM hallucinations through analyses of the EigenScores. In their paper, Chen et al. (2024) define a new metric for detecting confabulations, a subclass of hallucinations. They do this by calculating an EigenScore 3.4 based on determinant calculations from multiple outputs of an LLM with a high-temperature setting (temperature set to 0.5) to encourage the LLM to produce a variety of different outputs. They propose that if an LLM is set to hallucinate on that output, the generated texts will show higher semantic variability and produce a higher EigenScore. This method achieves SOTA performance and is unsupervised as it only relies on the representations learned by the model. In the forthcoming sections, we will analyze the correlation between the EigenScore of data points during training checkpoints and the most Sensitive Neurons associated with them.
@@
3.2 Sensitive Neuron Impact on EigenScores
To assess the correlation between Sensitive Neurons and other neurons in the embedding matrix of 10 generated outputs at a specific checkpoint, we conduct experiments aimed to determine if the presence of Sensitive Neurons indicates higher uncertainty and a greater likelihood of hallucinations.
We evaluate the Sensitive Neuron effect on the HELM dataset (Su et al., 2024), which includes outputs and internal states from six open-source LLMs based on inference over 50,000 Wikipedia articles, with human annotators labeling passages as factual or hallucinatory. To assess the impact of Sensitive Neurons on hallucination, we adapt the EigenScore method by applying it to sentence embeddings from the penultimate layer of EleutherAI’s Pythia 1B model, focusing on checkpoints between 133,000 and 143,000 training steps, where embeddings are more stable and the model has a higher degree of language understanding compared to initial checkpoints. We perform sensitive neuron dropout, removing the top 10% of Sensitive Neurons at each checkpoint, and compare the results to a baseline where 10% of neurons are randomly dropped. Additionally, we analyze the impact on hallucination-prone inputs versus non-hallucination-prone inputs to determine if Sensitive Neurons play a critical role during hallucination, without negatively affecting correct outputs.
3.2.1 What is the Effect of Sensitive Neurons on Hallucination Metrics?
Since a reduction in the EigenScore metric can be used as a proxy to show the reduction in likelihood of hallucination and an increase in the uncertainty of the model about its response, we keep using this metric in our investigations. We are able to show through our comparison of the baseline random neuron dropout and Sensitive Neuron dropout that Sensitive Neurons significantly reduce the EigenScore metric and in turn, reduce the possibility of a confabulation (Figure 2(a)) implying their highly important role in determining model’s certainty. Not only do we observe this in hallucinatory outputs, we also observe a smaller reduction in EigenScore when applying this technique to correctly answered queries (Figure 2(b)). This result indicates that our methodology has a significant effect on the uncertainty shown by an LLM. We observe that looking at the internal states of the model is an effective way to eliminate confabulating text generation in various model sizes.

### v2

# Hallucination Detox: Sensitivity Dropout ( SenD) for Large Language Model Training

@@ Abstract
As large language models (LLMs) are increasingly deployed across various industries, concerns regarding their reliability, particularly due to hallucinations—outputs that are factually inaccurate or irrelevant to user input—have grown. Our research investigates the relationship between the training process and the emergence of hallucinations to address a key gap in existing research that focuses primarily on post hoc detection and mitigation strategies. Using models from the Pythia suite (70M–12B parameters) and several hallucination detection metrics, we analyze hallucination trends throughout training and explore LLM internal dynamics. We introduce Sensitivity Dropout ( SenD), a novel training protocol designed to mitigate hallucinations by reducing variance during training. SenD achieves this by deterministically dropping embedding indices with significant variability, referred to as Sensitive Embedding Indices. In addition, we develop an unsupervised hallucination detection metric, Efficient EigenScore (EES), which approximates the traditional EigenScore in 2x speed. This efficient metric is integrated into our protocol, allowing SenD to be both computationally scalable and effective at reducing hallucinations. Our empirical evaluation demonstrates that our approach improves LLM reliability at test time by up to 40% compared to normal training while also providing an efficient method to improve factual accuracy when adapting LLMs to Wikipedia, Medical, and LegalBench domains.
@@ 1.1 Motivation
As Large Language Models (LLMs) become more sophisticated and widespread across industries, concerns about their reliability and safety have grown due to misuse and user errors. One of these concerning areas discovered by the scientific community is the phenomenon of hallucinations - LLMs producing content that may not align with real-world facts, the user’s input, or training data it has seen in the past (Huang et al., 2023a). In our research, we target a specific field of hallucinations called confabulations which occur when the LLM generates different responses given the same or similar inputs. This can be harmful when the generations alter between correct and factually incorrect responses.
@@
To explore these hallucination trends, we analyze models ranging from 70 million to 12 billion parameters within Pythia suite (Biderman et al., 2023), assessing them across various training checkpoints and tasks. Our goal is to validate the oscillatory behavior observed by Li et al. (2024) through evaluation metrics including HaluEval (Li et al., 2023), FactScore (Min et al., 2023), SelfCheckGPT (Manakul et al., 2023), and XSum (Narayan et al., 2018). Utilizing the reliability of internal model dynamics for quantifying hallucination likelihood, we use EigenScore (Chen et al., 2024) and Semantic Entropy (Kossen et al., 2024) to detect hallucination risk by analyzing variability in high-temperature outputs. Experiments utilize EigenScore and the HELM dataset (Su et al., 2024) to identify hallucinations during training.
In response to this variance, we introduce a novel training protocol called Sensitivity Dropout ( SenD). SenD is designed to emphasize confident learning of facts, and in turn reduce the likelihood of confabulations, rather than solely minimizing the training loss. By selectively dropping Sensitive Embedding Indices—those exhibiting significant fluctuations throughout training— SenD acts as a technique that reduces hallucination variance and enhances the model’s factual certainty. This provides a more reliable criterion for determining training termination, ensuring models not only achieve loss convergence but also display stable factual confidence. To maintain efficiency as model size and inference count increase, we propose the Efficient EigenScore (EES), a novel metric for hallucination detection. EES replaces EigenScore (Chen et al., 2024), the primary metric used in our experiments, offering a scalable solution with high correlation to EigenScore.
@@
Sensitivity Dropout ( SenD), a training-time method designed to reduce hallucination variance and increase model factual confidence during training.
@@ 1.2 Related Work
The majority of research on hallucinations in language models has focused on detecting and mitigating this phenomenon rather than explaining its underlying causes. Recent techniques can be categorized into two main approaches: those that rely on output probabilities at inference time (Manakul et al., 2023; Joshi et al., 2017; Li et al., 2023) and those that utilize internal representations or hidden layers of the model (Su et al., 2024; Chen et al., 2024; Kossen et al., 2024). While the former has demonstrated effectiveness, the latter offers deeper insights but often comes with computational trade-offs. Additionally, methods like Reinforcement Learning with Human Feedback (RLHF) have gained traction for enhancing model reliability (Yu et al., 2024). However, many of these post-hoc solutions enhance factual accuracy by layering algorithms atop pre-trained models, which can be inefficient. Our work addresses this gap by focusing on the internal dynamics of the model that contribute to hallucinations.
Regularization techniques have been introduced to fix the issue of variability, notably random neuron dropout, used to reduce the variance and ensure that no neuron is overpowering others (Srivastava et al., 2014; Baldi & Sadowski, 2013). Work such as that done by Santra et al. (2020); Ba & Frey (2013) aims to modify random neuron dropout to change the way neurons are dropped to a more deterministic, precise manner. This has allowed the authors to drop unimportant connections in a deep neural network to ensure that class discriminative information is propagated through the model correctly (Santra et al., 2020). Inspired by this, our aim is to target hallucinatory embedding indices in our models to ensure that factual information is propagated through. State-of-the-art hallucination metrics, especially those based on internal model dynamics, rely on spectral analysis and embedding matrix computations. Methods like EigenScore (Chen et al., 2024) and Semantic Entropy (Kossen et al., 2024) effectively assess hallucination risk but require multiple inferences, making them computationally demanding as models scale. Tools such as the Density of States (DOS) and the kernel polynomial method (KPM) have been explored to approximate spectral properties efficiently (Huang et al., 2023b; Lin et al., 2014). Building on these advancements, our work integrates efficient spectral analysis methods into hallucination detection, demonstrated through EES and SenD.
@@ 2 Oscillatory Behaviour Validation
The training checkpoints of a transformer model can be vital in understanding the dynamics of how the model learns. Beyond the model’s architecture and the data itself, numerous factors influence the learning process such as: whether the loss function penalizes the learner for factual mistakes it makes or if it primarily tries to force the model to memorize the data. While our paper does not aim to address the broader debate on whether LLMs truly understand language or rely on memorization, our analysis of training dynamics through multiple checkpoints shows that converging the training loss does not necessarily correspond to reducing hallucinations, verifying the results by Li et al. (2024) for LLM oscillatory hallucination behaviour during training. Our study utilizes Eleuther AI’s Pythia and LMEval tools (Biderman et al., 2023; Gao et al., 2024a) to examine the development and evolution of LLMs throughout the training process. Pythia comprises a suite of 16 LLMs, all trained on public data in the same sequential order, with sizes ranging from 70 million to 12 billion parameters. We use 20 equally spaced training checkpoints from the start to the finish for our analysis. These models are evaluated at each checkpoint on a variety of hallucination metrics such as SelfCheckGPT for Self-consistency (Manakul et al., 2023), XSum for Summarization (Narayan et al., 2018), Perplexity, and HaluEval (Li et al., 2023) for Question Answering (QA) tasks. High SelfCheckGPT Score means that the model is more likely to contradict itself on the given input, a higher Rouge1 score on XSum means the data is aligning better with the provided reference summary, lower perplexity implies higher prediction confidence, and higher Exact Match, Accuracy, and Correctness implies better performance of the model on QA tasks.
@@ 2.1 How do the established iterative training processes influence LLM hallucinations?
The analysis of hallucination oscillations, as shown in Figure 1, indicates a consistent pattern across different models: oscillations persist throughout training from the initial to the final checkpoint. This finding highlights the uncertainty of halting training solely based on the convergence of training loss. For instance, in QA settings, the optimal Exact Match of the outputs with ground truths is achieved in earlier checkpoints. This evidence challenges the notion that optimizing solely for unsupervised loss in SGD guarantees learning the most accurate representation of the data. This observation is seen more drastically in 1(a), where model size has nearly no effect on the performance of SelfCheckGPT. Instead, we observe oscillatory behaviour within self-consistency, implying that model size is not much effective at tackling the issue of confabulations verified by results in Appendix A.2 as well. Our mitigation approach ( SenD) is discussed in Section 4.
@@ 2.2 How does model complexity affect the emergence of hallucinations throughout training?
An analysis of hallucination detection metrics reveals a diminishing rate of improvement with increased model scaling, particularly up to the 12B parameter size (Appendix A.2 for the study). This suggests that beyond a certain point, even though there is improvement in the hallucinations, larger models do not significantly reduce hallucinations, indicating that scaling alone is not sufficient for building robust models. Instead, more refined approaches are needed to address the underlying variability in model behavior. For the following experiments, we focus on the Pythia 1B model.
@@ 3.1 Sensitive Embedding Indices
Where is the activation of one input , is the number of tokens in the sequence, is the token embedding activation matrix, and is the subtraction to get the penultimate layer index and the formula is adapted from Su et al. (2024). The penultimate layer of the LLM, being the layer closest to the output probabilities, is our primary focus for hallucination analysis due to its rich information about output certainty.
@@
Let denote the embedding of data point at embedding index of the contextual embedding after checkpoint . Then we define the net change formula as
With these definitions, we can now describe the crux of our investigation: Sensitive Embedding Indices (SEIs). These SEIs give us key parts of the model that we will prove contribute to the hallucination of LLM models. They can be used to adapt training procedures for lowering hallucination variation during training and better overall confidence at inference time. In essence, SEIs are embedding indices in the sentence embedding from definition 3.1 that experience drastic changes between checkpoints of the training, something we believe is related to the oscillatory behaviour in hallucination performance. When finding the most sensitive embedding indices, we typically want to select the top embedding indices for a specific data point’s representation. In our investigation we set .
Definition 3.3 ( Sensitive Embedding Indices - SEIs).
@@
where is the total variability during the last checkpoints and the most sensitive embedding indices are
@@
The above definition of SEIs is then applied to LLM hallucinations through analyses of the EigenScores. In their paper, Chen et al. (2024) define a new metric for detecting confabulations, a subclass of hallucinations. They do this by calculating an EigenScore 3.4 based on determinant calculations from multiple outputs of an LLM with a high-temperature setting (temperature set to 0.5) to encourage the LLM to produce a variety of different outputs. They propose that if an LLM is set to hallucinate on that output, the generated texts will show higher semantic variability and produce a higher EigenScore. This method achieves SOTA performance and is unsupervised as it only relies on the representations learned by the model. In the forthcoming sections, we will analyze the correlation between the EigenScore of data points during training checkpoints and the most sensitive embedding indices associated with them.
@@
3.2 Sensitive Embedding Index Impact on EigenScores
To assess the correlation between SEIs and other indices in the embedding matrix of 10 generated outputs at a specific checkpoint, we conduct experiments aimed to determine if the presence of SEIs indicates higher uncertainty and a greater likelihood of hallucinations.
We evaluate the SEI effect on the HELM dataset (Su et al., 2024), which includes outputs from six open-source LLMs based on inference over 50,000 Wikipedia articles, with human annotators labeling passages as factual or hallucinatory. This dataset was selected as Wikipedia is one of the main fact sources people refer to, therefore LLMs should be robust to this type of information as well as it being created with the intended use case of hallucination detection. To assess the impact of SEIs on hallucination, we adapt the EigenScore method by applying it to sentence embeddings from the penultimate layer of EleutherAI’s Pythia 1B model, focusing on checkpoints between 133,000 and 143,000 training steps, where embeddings are more stable and the model has a higher degree of language understanding compared to initial checkpoints. We perform SEI dropout, dropping the top 10% of SEIs at each checkpoint, and compare the results to a baseline where 10% of embedding indices are randomly dropped. Additionally, we analyze the impact on hallucination-prone inputs versus non-hallucination-prone inputs to determine if SEIs play a critical role during hallucination, without negatively affecting correct outputs.
3.2.1 What is the Effect of Sensitive Embedding Indices on Hallucination Metrics?
Since a reduction in the EigenScore metric can be used as a proxy to show the reduction in likelihood of hallucination, we keep using this metric in our investigations. We are able to show through our comparison of the baseline random embedding index dropout and SEI dropout that SEIs significantly reduce the EigenScore metric and in turn, reduce the possibility of a confabulation (Figure 2(a)), implying their highly important role in determining model’s certainty. Not only do we observe this in hallucinatory outputs, we also observe a smaller reduction in EigenScore when applying this technique to correctly answered queries (Figure 2(b)). This result indicates that our methodology has a significant effect on the uncertainty shown by an LLM. We observe that the internal states of the model are effective in the elimination of confabulating text generation in various model sizes.